Displaying reports 61-80 of 1084.Go to page 1 2 3 4 5 6 7 8 9 10 End
Reports until 12:35, Friday 18 April 2025
Vacuum Systems General (GQuEST)
Daniel Grass - posted 12:35, Friday 18 April 2025 - last comment - 14:44, Friday 18 April 2025(12253)
Nitronic 60 1/4-20 1/4 long helicoils and custom 1/4-100 Bushings Cleaned and are Baking Out

I cleaned 15 Nitronic 60 1/4-20 1/4 long helicoils and 4 of the 5 custom 1/4-100 bushings. I used the standard 1:30 simple Green:DI water, DI water, then isopropanol baths and sonication for 10 minutes each. I am currently baking them out at 200°C for 48 hours.

Comments related to this report
Daniel Grass - 14:44, Friday 18 April 2025 (12254)

200°C reached at 2:40 PM.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 23:25, Thursday 17 April 2025 - last comment - 17:03, Monday 21 April 2025(12251)
Stanford Research Systems Residual Gas Analyzer 100 Turned On and Data Taken; 1.2E-3 Pressure on LFC for unknown reason

[Jeff, Daniel]

I connected the Stanford Research Systems Residual Gas Analyzer 100 (SRS 100) RS232 cable (9 pin D-Sub) to a RS232 (male) to USB-A (male) adapter and plugged it into the USB hub in the mobile clean room. I also connected the same adapter from the turbo controller to the hub. I turned on the RGA and took some spectra around 3:30 today. I've attached the graphs from the data. The y-axis is Torr. The total pressure is the sum of the partial pressures; these graphs imply a pressure of around 1E-5 Torr, even though the Agilent pressure gauge had a reading around 1E-7 Torr, two orders of magnitude different. Jeff and I came in around 11 PM tonight and took another RGA reading; H2 and N2 levels were higher than at 3 PM, up to 3E-5 Torr. We then found out the pressure is 1E-3 Torr on the Agilent pressure gauge and the ion pump was off. I don't know if the ion pump going off is a cause or an effect of the high pressure. If the 6E-10 Torr/s rate of rise measurement is accurate, then the pressure after ~12 hours would be ~3E-5 Torr. This seems to indicate the high pressure caused the ion pump to go off. If the helium from the helium leak check were in the system, it would be easily visibile with the RGA (and shouldn't cause such a huge pressure rise). The up-to-air valve and the gate valve to the turbo pump were closed and tight.

If the pressure is this high, this explains the RGA measurments being so high (it's not rated above 1E-4 Torr, so a little discrepancy there makes sense). There is a big mystery for why the pressure is 1.2E-3 Torr after being stable for days. To test the agilent pressure gauge, I rotated and moved the magnet's z-axis (it's a cylinder and on another cylinder). This lowered the pressure to ~7E-4 Torr, half as low. A factor of 2 in pressure isn't a big deal for these issues.

The H2 levels on the RGA also seem quite high.

Images attached to this report
Comments related to this report
Daniel Grass - 11:26, Friday 18 April 2025 (12252)

Even though everything is off, the presssure went down to 2.8E-4. I don't have a good explanation. Maybe there are E&M fields from the RGA or ion pump that affect the pressure gauge's reading.

Daniel Grass - 17:03, Monday 21 April 2025 (12258)

See this post for a partial explanation for why the pressure rose to 1E-3 Torr and stabalized there.

Vacuum Systems General (RbQ)
Daniel Grass - posted 19:28, Wednesday 16 April 2025 (12250)
Infleqtion PICAS 2D Rb MOT Adapter Plate Attached to 2.75 in CF Bellows

[Jeff, Daniel]
We attached the Infleqtion PICAS 2D Rb MOT Adapter Plate to a 2.75" CF Bellows with 1.25" long 1/4-28 screws and no washers. 0.875" long screws weren't quite long enough. The gasket was initially incorrectly seated on the knife edges; putting the curved edge on the bellows seemed to help when I tried again with a new gasket. I tightened the bolts to 14 Nm, the standard for 2.75" CF thru flanges, and made steel to steel contact all around.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 16:51, Tuesday 15 April 2025 (12248)
In Vacuum Piezo Assembly Cleaned and Baking

I cleaned the 6082 aluminum parts that hold the mirror and piezo to lock in vacuum cavities/interferometers. I used the standard 1:30 Simple Green:DI water, DI water, then isopropanol baths and sonication for 3 minutes each. I am currently baking them out at 120°C for 48 hours.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:28, Monday 14 April 2025 - last comment - 18:31, Tuesday 15 April 2025(12245)
Laser Filter Cavity Vented and SRS 100 Residual Gas Analyzer Installed on Laser Filter Cavity

[Jeff, Ian, Daniel]

We vented the turbo/scroll system on the Laser Filter Cavity (LFC) by turning the knob on the turbo pump. There was a decent amount of air that came in, implying there is a good seal from the turbo to the outside world, even when the scroll and turbo are off. The gate valve was closed and we noticed no pressure increase in the main body of the LFC, implying the gate valve seal is good. We then vented the main body of the LFC with the up-to-air valve. A little turn increased the pressure from 1E-5 Torr to 1 Torr, so one needs to be careful there. I turned the valve a bit more and the pressure slowly increased. Notable, there was a lot of gas sucked into the LFC even when the pressure gauge read "over pressure" (over 760 Torr). I therefore tried to play with the magnet on the pressure gauge once the LFC was fully vented to get the controller to display 760 Torr. Rotating it and moving its z-position didn't accomplish it.

We took off a CF 2.75" to KF50 adapter + a KF50 blank flange and added a 7" long 2.75" CF tube to the all metal angle valve (the latter of which was quite dirty on the flange outside the vacuum ). We needed this 7" long part so that the cup of the Stanford Reserach Systems Residual Gas Analyzer 100 (SRS RGA 100) had room to be inserted into the vacuum system (it needs ~3"). We then connected the SRS RGA to the tube and we tightened the bolts.

I opened up the gate valves and had previously opened the all metal angle valve to the RGA and shut the up-to-air valve and turned on the scroll pump. After 20 minutes, the pressure was 0.1 Torr like before, so I don't think fiddling with the magnet changed anything. I then turned on the turbo (still in soft start mode) around 0.08 Torr. The pressure and power drawn behaved like last time. 20 minutes after turning on the turbo, the pressure is 4.9E-6 Torr.

To use the RGA (and have the turbo controller, which reads out the pressure) talk to the computer, we need some RS232 (male) to USB-A (male) adapers. Ian found this which looks good.

Images attached to this report
Comments related to this report
Daniel Grass - 22:07, Monday 14 April 2025 (12246)

The pressure was 9.8E-7 Torr at 7:42 PM.

Daniel Grass - 13:57, Tuesday 15 April 2025 (12247)

The pressure was 1.5E-7 Torr at ~1:40 PM. I turned on the ion pump and the pressure rose to 1E-5 Torr before dropping to 3E-7 Torr in a few minutes with the turbo still on. The pressure topped out around 3.9E-7 Torr before dropping to 3.8E-7 Torr.

Daniel Grass - 18:31, Tuesday 15 April 2025 (12249)

Yesterday, we had inserted the electronics control unit (ECU) onto the probe. I tightened the thumb screws to secure it. Since the pressure was 3.3E-7 Torr, I plugged in the ECU's power cord into the power strip. The green light turned on. I also installed the software onto the lab computer in the B111B control room. We don't yet have the RS232 to USB adapter to connect them, so I turned off the RGA and unplugged it.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:13, Monday 14 April 2025 (12244)
Rate of (Pressure) Rise Test for Laser Filter Cavity Performed

[Jeff, Daniel]

We performed a rate of rise test for the Laser Filter Cavity (LFC) to determine the outgassing/leak rate (it's hard to experimentally differentiate without a bakeout or measurment over time). I turned off the Ion Pump and the pressure rose from 2.2E-7 Torr to 1.5E-6 Torr in 3 minutes from the ion pump releasing some gas. The pressure then rose linearly at 6E-10 Torr/s. Assuming the volume is ~33 liters, this gives a leak rate + outgassing of 2.7E-8 mbar*L/s. The surface area is around 1 m^2 = 1E4 cm^2. Assuming the leak rate is negligible, the desorption rate is 3E-12 L*mbar/(cm^2*s), which is excellent (almost unbelivable). There are calculations that can be made from this data to access the performance of the vacuum pumps. I might do that later. See below for the raw data and graphs.

Time (minutes) Time (seconds) Pressure (Torr)
0 0 2.2E-07
0.666666667 40 7.5E-07
3 180 1.5E-06
3.5 210 1.6E-06
7.5 450 1.7E-06
8 480 1.8E-06
11 660 1.9E-06
18 1080 2.0E-06
20 1200 2.1E-06
24.5 1470 2.2E-06
32.5 1950 2.6E-06
35.7 2142 2.7E-06
38.25 2295 2.8E-06
Images attached to this report
Vacuum Systems General (GQuEST)
Daniel Grass - posted 23:54, Sunday 13 April 2025 (12243)
10 in to 6 in CF Zero Length Reducer Flange and 6 in CF 6 way cross added to IFO Cube

I cleaned the top of the IFO Vacuum Cube and added a 10" to 6" CF Zero Length Reducer Flange. I then added a 6" CF 6 way cross (with a rotatable flange face down). I tightened the bolts so that the copper flange wasn't visible and there was steel to steel contact all around.

I moved a bunch of vacuum equipment from B150 to B111B for the igo cube: A 2.75" CF Tee, 4 way cross, angle valve, all metal angle valve, pressure gauge (down to ~1 mTorr) two 6" CF Gate valves, a small ion pump, and an ion pump controller. A bunch of other equipment was attached to what we needed, including two 8" to 6" CF reducer tees, 6" CF tubes, additional 2.75" angle valves.

I also moved an all metal angle valve into B111A for the RbQ chamber.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 23:14, Friday 11 April 2025 - last comment - 23:38, Sunday 13 April 2025(12239)
Laser Filter Cavity Ion Pump Turned On

[Lee, Jeff, Daniel]

We plugged in the ion pump controller to a wall outlet and read the current from the wall outlet to the ion pump controller as 0.03 A. The voltage from the BNC on the front of the controller was 9 V.

The pressure was 2.5E-7. We hit enable and the on button around 2:40 PM. The pressure jumped to 1E-4 Torr and then quickly dropped. The BNC voltage dropped to 2 V before slowly rising to around 3.1 V. The controller drew 0.7-0.9 A. There was no need to repeatedly hit the on button; the ion pump stayed on and did not trip.

At 4:28, the pressure was 1.9E-7 Torr. I shut the gave vale to the turbo pump and the pressure rose to 3.3E-7 Torr. I turned off the turbo pump and the pressure again jumped to 3.8E-7 Torr. By 4:47 PM, the pressure was 4.7E-7 Torr. No turbo pump speed was given and it sounded quiet, so I turned off the scroll at 4:49 PM. The BNC voltage was 3.95 V.

By 4:54 PM, the pressure was 5.1E-7 Torr and the BNC voltage was 3.93 V.

After turning the turbo off, I believe I enabled a setting so that the turbo speed will be displayed as its spinning down. The controller is unfortunately hard to use.

I will monitor the pressure in the coming days. If the pressure approaches 1E-4 Torr, I am going to turn off the ion pump to protect it.

Going forward, it might be nice to add a valve to the ion pump so that all systems can be vented nicely. Maybe one of these up-to-air valves (cheaper than on Lesker) and a CF Tee.

Images attached to this report
Comments related to this report
Daniel Grass - 23:11, Friday 11 April 2025 (12241)

At 11 PM, the pressure went down to 4.4E-7 Torr. The BNC voltage was 4.09 V and the drawn current from the wall was 0.587A. The fact that the pressure dropped with only the ion pump on should imply the ion pump actually works.

Daniel Grass - 23:38, Sunday 13 April 2025 (12242)
Time Pressure (Torr) BNC Voltage (V)
12:37 AM Saturday 4.2E-7 Not measured
5:31 PM Saturday 3.1E-7 4.29
7:52 PM Saturday 3.0E-7 4.32
7:29 PM Sunday 2.5E-7 4.44

 

I want to do a rate of rise test to evaluate the ion pump and turbo's performance, assuming outgassing (which decreases over time) is neglibigle.

Lab Infrastructure General
Alex Ramirez - posted 19:15, Friday 11 April 2025 (12240)
Google Doorbell Update

I have successfully gotten the Google doorbell to work via python

However, Google restricts your ability to pull media from the doorbell api if you are not using their home app, so when I see the video clips, it returns a 401 error, which indicates that I am not accessing via the home app

There is a work around, but this will take some time to test and may not be a priority: https://medium.com/@tamirmayer/google-nest-camera-internal-api-fdf9dc3ce167

My code currently uses the Google API and a refresh token to try to access the image preview (we will see if it works after trying the thing above)

The doorbell now uses pub/sub to post events, and my python code allows the user to subscribe to the event log and pull event threads when available. The events for a doorbell press (sdm.devices.events.DoorbellChime.Chime) require no other user interaction and can trigger the doorbell notification via Mattermost once I work with Ian to set this up. The result is:
 

Listening for messages on projects/lab-doorbellapi/subscriptions/doorbell-subscription-v1...

Received message: b'{\n  "eventId": "b3326428-aceb-461a-ab2a-25ab3f8c06a6",\n  "timestamp": "2025-04-12T02:00:15.064852Z",\n  "resourceUpdate": {\n    "name": "<redacted>,\n    "events": {\n      "sdm.devices.events.DoorbellChime.Chime": {\n        "eventSessionId": "1740690860",\n        "eventId": "n:1"\n      }\n    }\n  },\n  "userId": "<redacted>",\n  "eventThreadId": "6855bbc9-f284-4582-a679-b875450c9ec2",\n  "resourceGroup": ["<redacted>"],\n  "eventThreadState": "STARTED"\n}'
Processing event for eventThreadId: 6855bbc9-f284-4582-a679-b875450c9ec2
Doorbell was pressed! 
Received message: b'{\n  "eventId": "b3326428-aceb-461a-ab2a-25ab3f8c06a6",\n  "timestamp": "2025-04-12T02:00:15.064852Z",\n  "resourceUpdate": {\n    "name": "<redacted>",\n    "events": {\n      "sdm.devices.events.DoorbellChime.Chime": {\n        "eventSessionId": "1740690860",\n        "eventId": "n:1"\n      }\n    }\n  },\n  "userId": "<redacted>",\n  "eventThreadId": "6855bbc9-f284-4582-a679-b875450c9ec2",\n  "resourceGroup": ["<redacted>"],\n  "eventThreadState": "ENDED"\n}'
Processing event for eventThreadId: 6855bbc9-f284-4582-a679-b875450c9ec2
 

The print statement "Doorbell was pressed!" occurs when a new event is processed and its eventThreadState is "STARTED" 

This will be changed to a Mattermost post via webhook in the future

Access to the developer settings for the device console, pub/sub, API, and Google Cloud developer settings are all set up for gquestlab@gmail.com 


Please don't make any changes to the Lab-DoorbellAPI project without contacting me first.

 

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:10, Thursday 10 April 2025 - last comment - 11:51, Friday 11 April 2025(12237)
Laser Filter Cavity Pumps turned on

[Ian, Jeff, Daniel]

We connected the IDP 7 Scroll Pump for the Laser Filter Cavity (LFC) to the same power system as the SNSPD Dewer pumps. We connected the Turbo and Turbo Controller to a power strip on the top shelf in the mobile clean rooms. The pressure gauge, after warming up for a few minutes, read "pressure too high". This is definitely a measurement error, but I think it's fine. I plugged in the USB-B extension cords from amazon and they successfully power the fan that cools the turbo pump.

We turned on the scroll pump at 4:24 PM, and after ~30 seconds it made a secondary "turning on" mechanical sound and the pressure gauge started to read actual values instead of "pressure too high". After 20 minutes, the pressure was 0.1 Torr. We plugged in the interlock bypass D-Sub and turned on the turbo. After 2 minutes (and 1 vertical bar on the turbo controller), the pressure was 3*10^-4, after 12 minutes, the pressure was 8.7*10^-6 Torr. The 3 vertical bars then revealed the turbo speed, 1167 Hz. The final power was 12 W after peaking around 90 W while ramping up. I left after 15 total minutes of the turbo being on with a pressure of 6.9*10^-6 Torr.

Comments related to this report
Daniel Grass - 11:51, Friday 11 April 2025 (12238)

Pressure at 6:53 PM last night: 1.2E-6 Torr

Pressure at 11:30 AM today: 2.8E-7 Torr

Vacuum Systems General (GQuEST)
Daniel Grass - posted 21:12, Wednesday 09 April 2025 (12235)
Vacuum Tubing Connected from Laser Filter Cavity Turbo Pump to Scroll

I connected the ~2 ft long KF 16 to KF 25 tubing from the Agilent TwisTorr 74 FS Turbo Pump to a 4 ft long KF 25 tubing which was connected to the Agilent IDP 7 scroll pump. I used the included filter inside the o-ring going to the scroll pump.

Now that the vacuum system is fully connected, I opened up the gate valves to the turbo pump and the ion pump. I also connected the power cord to the scroll pump but didn't plug it into a wall. I flipped a red switch so that the scroll pump is expecting 110-125 V instead of 230 V.

Images attached to this report
Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:25, Wednesday 09 April 2025 - last comment - 23:09, Wednesday 09 April 2025(12234)
SNSPD Dewer Fully Assembled Pumped out with Agilent IDP 7 Scroll Pump and TwisTorr 74 Turbo Pump

[Alex, Daniel]

Alex finished the assembly of the SNSPD Dewer and vacuum parts. We then pumped out the vacuum system with the IDP 7 scroll pump to 8.7*10^-2 Torr (there is some systematic uncertainty with the pressure measurements with the pressure gauge and the magnet location). We enabled soft start on the turbo pump rack controller (see the wiki for how to) and plugged in the pump (while the rack controller wasn't powered). The soft start, which slowly ramps up the pump speed which is good when the pump has been off for over 3 months) remained on after this power cycle. We plugged in the included female 9 pin D-Sub with pins 3 and 8 shorted into P1 to bypass the interlock. We could then turn on the turbo pump. The pressure immediately dropped to ~10^-3 Torr and the turbo increased up its speed (not displayed by the controller) over ~15 minutes. The power draw by the turbo started at 10 W and peaked around 90 W. After these 15 minutes, the pressure was 1.6*10^-4 Torr and the rotation speed was 1167 Hz, the nominal value. The power drawn was 16 W.

Comments related to this report
Daniel Grass - 23:09, Wednesday 09 April 2025 (12236)

Pressure values later on:

Turbo on around 5 PM

Pressure at 7:30 PM: 6.7*10^-5 Torr

Pressure at 8 PM: 6.4*10^-5 Torr

Pressure at 10:38 PM: 5.7*10^-5 Torr

Displaying reports 61-80 of 1084.Go to page 1 2 3 4 5 6 7 8 9 10 End